Question3 Expand the given expression. Type th 4 In A) 12In(x +1)+ 20 In( 1) -2ln( B) 4 ln
Answer #1

Given expression is = \ln\left\(\frac{(x+1)^3(x-1)^5}{\sqrt{2x^2-x-1}}\right\)^4

= 4\ln\left\(\frac{(x+1)^3(x-1)^5}{\sqrt{2x^2-x-1}}\right\)

= 4[\ln\left\((x+1)^3(x-1)^5\right\)-\ln(\sqrt{2x^2-x-1})]

= 4[\left\{ \ln\left\((x+1)^3\right\)+\ln((x-1)^5) \right\}-\ln(2x^2-x-1)^{1/2}]

= 4[3\ln(x+1)+5\ln(x-1)-\frac{1}{2}\ln(2x^2-x-1)]

= 12\ln(x+1)+20\ln(x-1)-\frac{4}{2}\ln(2x^2-x-1)

= 12\ln(x+1)+20\ln(x-1)-2\ln(2x^2-x-1)